OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (-2, 1). Which equation defines circle P?

hartnn (hartnn):

a general equation of circle is \((x-h)^2+(y-k)^2=r^2\) where h,k is centre r is radius.

OpenStudy (anonymous):

\[(x+2)^{2}+(y-1)^{2}=(2\sqrt{2})^{2}\]

hartnn (hartnn):

=8^2

hartnn (hartnn):

u get that @Patastic81 ?

OpenStudy (anonymous):

sorry it will be \[8^{2}\]

hartnn (hartnn):

\((x+2)^{2}+(y-1)^{2}=8^2\)

OpenStudy (anonymous):

@Patastic81 it should be 8^2

hartnn (hartnn):

there are 3 tabls below Equaion Draw Attach File did u see those ?

hartnn (hartnn):

*Equation

hartnn (hartnn):

u click on Attach Files, and u can upload the pic.

hartnn (hartnn):

when u click on Attach Files. a new window will appear.

hartnn (hartnn):

where are those points , are those random or they form diameter ?

hartnn (hartnn):

after u attached, u click Post.

hartnn (hartnn):

means u have pic in your computer, right ?

hartnn (hartnn):

so where is the pic ? u can take screenshot and save it in your computer.

hartnn (hartnn):

'the pic is on my comp... but its not saved' i am not sure what that means ....

hartnn (hartnn):

can u take screenshot ? i don't think u can save the assignment in your computer...

hartnn (hartnn):

thats clear

hartnn (hartnn):

first of all, u can observe that the centre is in 2nd quadrant, right ?

hartnn (hartnn):

so, centre is ...?

hartnn (hartnn):



hartnn (hartnn):

radius is length of segment from centre to any point on circle.

hartnn (hartnn):

new question, new post please :)

hartnn (hartnn):

see my first comment, and tell me the centre.

hartnn (hartnn):

h=.... ? k=.....?

hartnn (hartnn):

before attaching circles, please find centre.

hartnn (hartnn):

see my first comment. compare your equation with that equation and tell me what u get for h=....? k=......? r=...... ?

hartnn (hartnn):

i can only delete my comments.

hartnn (hartnn):

thats why we say, new question , new post.....but lets do this one here.

hartnn (hartnn):

u had your equation, compare it with this eq. \((x-h)^2+(y-k)^2=r^2\)

hartnn (hartnn):

and find, h,k,r

OpenStudy (anonymous):

let me post all four pics first then we start okay? because it wont allow me to post all four at same time okay?

hartnn (hartnn):

we can do this without options..

OpenStudy (anonymous):

my answer is a picture, please just let me post?

OpenStudy (anonymous):

first one

hartnn (hartnn):

just give the equation.

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

third one

OpenStudy (anonymous):

fourth one

OpenStudy (anonymous):

those are my 4 possible answers

OpenStudy (anonymous):

the questions is: Circle T has an equation of (x + 10)2 + y2 = 25. Which graph represents circle T?

OpenStudy (anonymous):

how can i answer this question sir?

hartnn (hartnn):